OpenStudy (anonymous):

Evaluate x – 3y − 3z when x = −4, y = 1, and z = −1.

OpenStudy (ajprincess):

Substitute the values. x – 3y − 3z =-4-3(1)-3(-1) Can u do nw?

OpenStudy (anonymous):

yes. but when i substitute them, do i multiply?

OpenStudy (ajprincess):

x – 3y − 3z =-4-3*1-3*(-1) Nw can u do?

OpenStudy (anonymous):

yes. thanks ☺

OpenStudy (ajprincess):

yw.:)